14. What is the velocity of an electron with a de Broglie wavelength of 4.32 mm? 168 m/s 211 m/s O 300,000 m/s O 145 m/s
Answer #1

De Broglie's Equation :  λ = h/(m*v)

Where ,  λ = wavelength in meters

m = mass in kilograms

v = velocity in meter/second

h = Plank's constant = 6.626 * 10-34 J/Hz

Soluz De Broglies Equation a=h mv M= mass of of eletton = 9.11 X 10-3 1 Kg 7= 4.32mm = 4,32X10-3, h: 6.626X1D-34 J/Hz wo 7=h